OpenStudy (idealist10):

Solve (3x^2*cosxy-x^3*y*sinxy+4x)dx+(8y-x^4*sinxy)dy=0.

zepdrix (zepdrix):

\[\Large\rm (3x^2 cosxy-x^3y sinxy+4x)dx+(8y-x^4 sinxy)dy=0\]Is it exact? :o

OpenStudy (idealist10):

Integral of (8y-x^4*sinxy)dy=4y^2-y*x^4*sinxy+h(x)=C 0+h'(x)=3x^2*cosxy-x^3*y*sinxy+4x h(x)=x^3*cosxy-x^4/4*y*sinxy+2x^2 But the answer in the book is x^3*cosxy+4y^2+2x^2=C. Yes, it's an exact equation.

OpenStudy (idealist10):

I got x^3*cosxy-y*x^4*sinxy-x^4/4*y*sinxy+4y^2+2x^2=C. Which answer is right?

zepdrix (zepdrix):

Ugh it's so hard to read the text :(

zepdrix (zepdrix):

\[\Large\rm \int\limits 8y-x^4\sin xy~dy=4y^2+x^3\cos xy+ h(x)\] Hmm yours looks quite different from this..

OpenStudy (idealist10):

How did you get x^3*cosxy?

zepdrix (zepdrix):

\[\Large\rm \int\limits -\sin(ay)dy=\frac{1}{a}\cos(ay)\]

zepdrix (zepdrix):

We lose a factor of x out front when we integrate in y.

zepdrix (zepdrix):

Your sine function didn't change when you integrated :o that was strange hehe

zepdrix (zepdrix):

Looks like you applied the power rule to it by mistake :3

zepdrix (zepdrix):

Uh oh, did your brain explode? :( Ideal, noooooooooo!

OpenStudy (idealist10):

Then how come my Ti-89 calculator says the integral of -x^4*sinxy dy=-y*x^4*sinxy?

zepdrix (zepdrix):

Hmm, I'm not sure :o Calculators can be bonkers sometimes. Wolfram appears to be giving us the correct information though: https://www.wolframalpha.com/input/?i=integral+of+-x%5E4sin%28xy%29dy

zepdrix (zepdrix):

Stiill don't believe it? :o lol

OpenStudy (idealist10):

I do believe it. Right now, I'm working this problem out and see if it really makes sense. Please wait.

OpenStudy (idealist10):

Okay, what's the integral of -x^3*y*sinxy?

zepdrix (zepdrix):

with respect to which variable?

OpenStudy (idealist10):

dx

zepdrix (zepdrix):

I think you want to avoid integrating in x, that's gonna be so much by parts :( painful painful painful.

zepdrix (zepdrix):

\[\Large\rm \int\limits\limits 8y-x^4\sin xy~dy=4y^2+x^3\cos xy+ h(x)\]Taking partial with respect to x gives:\[\Large\rm 0+3x^2\cos xy-x^3y \sin xy+h'(x)\]And then we compare that to our thing with the dx on it.

OpenStudy (idealist10):

I got it wrong again while doing partial with respect to x.

zepdrix (zepdrix):

Taking our partial, respect to x: \[\Large\rm (4y^2)_x+(x^3 \cos xy)_x+h(x)_x\] Setting up product rule:\[\Large\rm =0+(x^3)_x \cos xy+ x^3(\cos xy)_x+h'(x)\]

zepdrix (zepdrix):

\[\Large\rm =0+(3x^2) \cos xy+ x^3(-y sinxy)+h'(x)\]

zepdrix (zepdrix):

Forget your product rule maybe? Differentiating that cosine can be a little tough also :)

OpenStudy (idealist10):

I got it! Thanks a lot for the help on this problem!

zepdrix (zepdrix):

yay team \c:/